### v1

# Small Hurewicz and Menger sets which have large continuous images

Abstract.
We provide new techniques to construct sets of reals without perfect subsets and with the Hurewicz or Menger covering properties. In particular, we show that if the Continuum Hypothesis holds, then there are such sets which can be mapped continuously onto the Cantor space. These results allow to separate the properties of Menger and ${\mathsf{S}_{1}(\Gamma,\mathrm{O})}$ in the realm of sets of reals without perfect subsets and solve a problem of Nowik and Tsaban concerning perfectly meager subsets in the transitive sense. We present also some other applications of the mentioned above methods.
Key words and phrases:
Hurewicz, Menger, totally imperfect, perfectly meager, everywhere meager, countably perfectly meager
2010 Mathematics Subject Classification:
Primary: 54D20
The research of the first and second authors was funded by the National Science Center, Poland Weave-UNISONO call in the Weave programme Project: Set-theoretic aspects of topological selections 2021/03/Y/ST1/00122. The research of the third author was funded in whole by the Austrian Science Fund (FWF) [I 5930].
1. Introduction
Throughout this paper, we assume that a set $X$ with a given topological property is a subset of the Cantor space ${2^{\omega}}$ or its square. A set $X$ is Menger if for any sequence $\mathcal{U}_{0},\allowbreak\mathcal{U}_{1},\allowbreak\dotsc$ of open covers of $X$, there are finite sets $\mathcal{F}_{0}\subseteq\mathcal{U}_{0},\mathcal{F}_{1}\subseteq\mathcal{U}_{1},\dotsc$ such that the family $\{\,\bigcup\mathcal{F}_{n}:n\in\omega\,\}$ covers $X$. If in addition the family $\{\,\bigcup\mathcal{F}_{n}:n\in\omega\,\}$ is a $\gamma$-cover of $X$, i.e., for every $x\in X$, the set $\{\,n:x\in\bigcup\mathcal{F}_{n}\,\}$ is co-finite, then the set $X$ is Hurewicz. Each $\sigma$-compact set is Hurewicz and each Hurewicz set is Menger. By the results of Bartoszyński–Shelah [5] and Tsaban–Zdomskyy [25], there are uniform constructions of sets in ZFC which show that none of these implications are reversible in the realm of sets of reals. The above properties are central in the combinatorial covering propeties theory and found applications in such areas as forcing [6], function spaces [20], Ramsey theory in algebra [24], combinatorics of discrete subspaces [1], hyperspaces with the Vietoris topology [9], products of Lindelöf spaces [2, 22] and products of paracompact spaces [21].
Each of the considered properties has a combinatorial characterization using continuous mappings into the Baire space ${\omega^{\omega}}$. A set $X$ is Menger (Hurewicz) if and only if every continuous image of $X$ into ${\omega^{\omega}}$ is non-dominating (bounded) [18]. It follows that each of these properties has a critical cardinality, i.e., the minimal cardinality of a subset of ${\omega^{\omega}}$ not having a given property. The critical cardinality for Menger is the dominating number $\mathfrak{d}$ and for Hurewicz it is the bounding number $\mathfrak{b}$. A set is totally imperfect if it does not contain a perfect subset, i.e., a homeomorphic copy of ${2^{\omega}}$. Known constructions of totally imperfect sets which are Menger or Hurewicz involve category-theoretic, measure-theoretic, combinatorial or forcing methods. However, each of these constructions provide a set satisfying ${\mathsf{S}_{1}(\Gamma,\mathrm{O})}$, which is another classical covering property well studied in the field [7, 19, 11]. A set $X$ is ${\mathsf{S}_{1}(\Gamma,\mathrm{O})}$ if for every sequence $\mathcal{U}_{0},\allowbreak\mathcal{U}_{1},\allowbreak\dotsc$ of open $\gamma$-covers of $X$, there are sets $U_{0}\in\mathcal{U}_{0},U_{1}\in\mathcal{U}_{1},\dotsc$ such that the family $\{\,U_{n}:n\in\omega\,\}$ covers $X$. The critical cardinality for ${\mathsf{S}_{1}(\Gamma,\mathrm{O})}$ is $\mathfrak{d}$ and the following implications hold [7].
[…]
The only ZFC examples for separation Hurewicz or Menger from ${\mathsf{S}_{1}(\Gamma,\mathrm{O})}$ are not satisfactory, i.e., they are the Cantor space and, more generally, sets containing perfect subsets. Prior to this research, there hadn’t been totally imperfect examples, even under some additional assumptions beyond ZFC.
Using various set-theoretic assumptions which are weak portions of the Continuum Hypothesis, we construct totally imperfect Hurewicz or Menger sets which can be mapped continuously onto the Cantor cube. Each of already known strategies to construct totally imperfect sets with these properties have substantial limitations which prevent from reaching such a goal. These results have consequences for separation of considered properties. Another benefit is a solution of the below problem which was formulated independently by Nowik and Tsaban (in oral communication). By $\oplus$ we mean a standard group operation in ${2^{\omega}}$, i.e., for every $x,y\in{2^{\omega}}$ we have
$(x\oplus y)(n):=(x(n)+y(n))(\bmod 2)$
for all natural numbers $n$. For a set $A\subseteq{2^{\omega}}$ and an element $t\in{2^{\omega}}$, let $t\oplus A:=\{\,t\oplus a:a\in A\,\}$. A set $X\subseteq{2^{\omega}}$ is perfectly meager if for any perfect set $P\subseteq{2^{\omega}}$, the set $X\cap P$ is meager in the relative topology of $P$. By the result of Just–Miller–Scheepers–Szeptycki [7, Theorem 5.5] any totally imperfect Hurewicz subset of ${2^{\omega}}$ is perfectly meager. A set $X\subseteq{2^{\omega}}$ is perfectly meager in the transitive sense [12, 14] (PMT in short) if for any perfect set $P\subseteq{2^{\omega}}$, there is a $\sigma$-compact set $F\subseteq{2^{\omega}}$ containing $X$ such that the set $F\cap(t\oplus P)$ is meager in the relative topology of $t\oplus P$ for all $t\in{2^{\omega}}$.
Theorem 1.1 (Nowik [13, Theorem 1.]).
Each Hurewicz set in ${2^{\omega}}$ which cannot be mapped continuously onto ${2^{\omega}}$ is PMT.
In view of Zakrzewski’s result [15] which states that every totally imperfect Hurewicz set is universally meager (see definition in Subsection 4.2), the following question seems to be natural.
Problem 1.2 (Nowik, Tsaban).
Is any Hurewicz totally imperfect subset of ${2^{\omega}}$, PMT?
We show that a solution to this problem is independent from ZFC.
2. Hurewicz sets
Let $\operatorname{cov}(\mathcal{N})$ be the minimal cardinality of a family of Lebesgue-null subsets of ${2^{\omega}}$ which covers ${2^{\omega}}$. A subset $X$ of a topological space is a $\lambda^{\prime}$-set if for any countable subset $A$ of the space the set $A$ is a $G_{\delta}$-set in the relative topology of $X\cup A$.
Theorem 2.1.
Assume that $\operatorname{cov}(\mathcal{N})=\mathfrak{b}=\mathfrak{c}$. Then there exists a set $X\subseteq{2^{\omega}}\times{2^{\omega}}$ with the following properties:
  1. (1)
    $X$ is Hurewicz,
  2. (2)
    $X$ is a $\lambda^{\prime}$-set,
  3. (3)
    $X$ is totally imperfect,
  4. (4)
    $\pi[X]={2^{\omega}}$,
  5. (5)
    a homeomorphic copy of $X$ in ${2^{\omega}}$ is not PMT.
Remark 2.2.
By the equivalence (1) $\Leftrightarrow$ (3) in Proposition 2.6 from [29], in the Laver model every $\lambda^{\prime}$-set is Hurewicz, and by the discussion before Lemma 2.3 in [17] we have that every Hurewicz $\lambda$-set in this model has size $\omega_{1}$. Since $\mathfrak{b}=\mathfrak{c}=\omega_{2}$ in the Laver model, we conclude that $\mathfrak{b}=\mathfrak{c}$ is not sufficient to achieve the disjunction of (2) and (4) in Theorem 2.1.
In order to prove Theorem 2.1 we need the following auxiliary results.
Lemma 2.3 ([14, Lemma 6]).
A set $X\subseteq{2^{\omega}}$ is PMT if and only if for any sequence $P_{0},\allowbreak P_{1},\allowbreak\dotsc$ of perfect sets in ${2^{\omega}}$, there are compact sets $F_{0},\allowbreak F_{1},\allowbreak\dotsc\subseteq{2^{\omega}}$ such that $X\subseteq\bigcup_{n\in\omega}F_{n}$ and $t\oplus P_{n}\not\subseteq F_{m}$ for every natural numbers $n,m$ and element $t\in{2^{\omega}}$.
For every set $A\subseteq{2^{\omega}}\times{2^{\omega}}$ and element $x\in 2^{\omega}$, let
$A|_{x}:=\{\,y\in{2^{\omega}}:(x,y)\in A\,\}.$
Let $\pi\colon{2^{\omega}}\times{2^{\omega}}\to{2^{\omega}}$ be the projection onto the first coordinate.
Lemma 2.4.
Let $A\subseteq{2^{\omega}}\times{2^{\omega}}$ be a clopen set such that $\pi[A]={2^{\omega}}$ and $M\subseteq{2^{\omega}}\times{2^{\omega}}$ be a closed set such that the sets $M|_{x}$ are nowhere dense in ${2^{\omega}}$ for all elements $x\in{2^{\omega}}$. Then there is a clopen set $B\subseteq A\setminus M$ such that $\pi[B]={2^{\omega}}$.
Proof.
Any set $A$ with the given property can be replaced by a finite union of nonempty basic open sets in ${2^{\omega}}\times{2^{\omega}}$. Since the sets $M|_{x}$ are nowhere dense in ${2^{\omega}}$ for all $x\in{2^{\omega}}$, it is enough to prove the statement for $A={2^{\omega}}\times{2^{\omega}}$. For each element $x\in{2^{\omega}}$ there are an element $y\in{2^{\omega}}$ and clopen basic sets $U_{x},V_{x}\subseteq{2^{\omega}}$ such that
$(x,y)\in U_{x}\times V_{x}\subseteq({2^{\omega}}\times{2^{\omega}})\setminus M.$
Then there is a finite set $X^{\prime}\subseteq X$ such that the family $\{\,U_{x}:x\in X^{\prime}\,\}$ covers ${2^{\omega}}$. Put
$B:=\bigcup\{\,U_{x}\times V_{x}:x\in X^{\prime}\,\}.\qed$
Lemma 2.5.
Let $M_{0},\allowbreak M_{1},\allowbreak\dotsc$ be an increasing sequence of closed sets in ${2^{\omega}}\times{2^{\omega}}$ such that the sets $M_{n}|_{x}$ are nowhere dense in ${2^{\omega}}$ for all natural numbers $n$ and every $x\in{2^{\omega}}$. Then there is a compact set $D\subseteq{2^{\omega}}\times{2^{\omega}}$ disjoint from $\bigcup_{n\in\omega}M_{n}$ such that $\pi[D]={2^{\omega}}$.
Proof.
Let $D_{0}:={2^{\omega}}\times{2^{\omega}}$. Fix a natural number $n$ and assume that a clopen set $D_{n}\subseteq{2^{\omega}}\times{2^{\omega}}$ with $D_{n}\cap M_{n}=\emptyset$ and $\pi[D_{n}]={2^{\omega}}$, has already been defined. Let $D_{n+1}$ be a set from Lemma 2.4, applied to the sets $D_{n}$ and $M_{n}$. Then the set $D:=\bigcap_{n\in\omega}D_{n}$ is disjoint from $\bigcup_{n\in\omega}M_{n}$ and $\pi[D]={2^{\omega}}$. ∎
Lemma 2.6.
Let $X\subseteq{2^{\omega}}\times{2^{\omega}}$ be a set such that for each open set $G\subseteq{2^{\omega}}\times{2^{\omega}}$ containing $X$ there is an element $x\in{2^{\omega}}$ with $G|_{x}={2^{\omega}}$. Then there is a homeomorphic copy of $X$ in ${2^{\omega}}$ which is not PMT.
Proof.
Let $K_{0},K_{1}\subseteq{2^{\omega}}$ be perfect sets such that the map $\oplus\colon K_{0}\times K_{1}\to{2^{\omega}}$ is injective and let $\phi_{0},\phi_{1}$ be homeomorphisms from ${2^{\omega}}$ onto $K_{0}$ and $K_{1}$, respectively. Then the map
$\Phi\colon{2^{\omega}}\times{2^{\omega}}\to K_{0}\oplus K_{1}$
defined by
$\Phi(x,y):=\phi_{0}(x)\oplus\phi_{1}(y),$
for all $(x,y)\in{2^{\omega}}\times{2^{\omega}}$, is a homeomorphism.
Assume that the set $\Phi[X]$ is PMT. Let $\{\,P_{n}:n\in\omega\,\}$ be an enumeration of a clopen basis of $K_{1}$. Applying Lemma 2.3 to the sequence $P_{0},\allowbreak P_{1},\allowbreak\dotsc$ and to the set $\Phi[X]$, there are compact sets $F_{0},\allowbreak F_{1},\allowbreak\dotsc\subseteq{2^{\omega}}$ such that $\Phi[X]\subseteq\bigcup_{n\in\omega}F_{n}$ and $t\oplus P_{n}\not\subseteq F_{m}$ for every natural numbers $n,m$ and elements $t\in{2^{\omega}}$. Consequently, the sets $F_{n}$ are nowhere dense in $t\oplus K_{1}$ for all $t\in{2^{\omega}}$. Let $M_{n}:=\Phi^{-1}[\bigcup_{i\leq n}F_{i}]$ for all natural numbers $n$. Then $X\subseteq\bigcup_{n\in\omega}M_{n}$ and the sets $M_{n}|_{x}$ are nowhere dense in ${2^{\omega}}$ for every natural number $n$ and each $x\in{2^{\omega}}$. Let $D\subseteq{2^{\omega}}\times{2^{\omega}}$ be a set obtained from Lemma 2.5. Then the open set $O:=({2^{\omega}}\times{2^{\omega}})\setminus A$ contains $X$ and $O|_{x}\neq{2^{\omega}}$ for all elements $x\in{2^{\omega}}$, which is a contradiction. Thus, the set $\Phi[X]$ is not PMT. ∎
Let $\mu$ be the Lebesgue measure on ${2^{\omega}}$.
Lemma 2.7.
Let $O\subseteq{2^{\omega}}\times{2^{\omega}}$ be an open set such that $\mu(O|_{x})=1$ for all elements $x\in{2^{\omega}}$. Then for every real number $a\in(0,1)$ there exists a compact set $K\subseteq O$ such that $\mu(K|_{x})>1-a$ for all elements $x\in{2^{\omega}}$.
Proof.
For each $x\in{2^{\omega}}$ there is an an open set $U_{x}\subseteq{2^{\omega}}$ such that ${2^{\omega}}\setminus O|_{x}\subseteq U_{x}$ and $\mu(U_{x})<a$. Then there is a clopen neighborhood $V_{x}\subseteq{2^{\omega}}$ of $x$ such that ${2^{\omega}}\setminus O|_{x^{\prime}}\subseteq U_{x}$ for all $x^{\prime}\in V_{x}$. Let $\{V_{x_{i}}:i<n\}$ be a finite subcover of the cover $\{V_{x}:x\in 2^{\omega}\}$ of $2^{\omega}$. Replacing $V_{x_{i}}$ with $V_{x_{i}}\setminus\bigcup_{j<i}V_{x_{j}}$, if necessary, we may assume in addition that the sets $V_{x_{i}}$ are disjoint. Define
$W:=\bigcup_{i<n}V_{x_{i}}\times U_{x_{i}}.$
Then
$({2^{\omega}}\times{2^{\omega}})\setminus O\subseteq W\text{\quad and\quad}\mu(W|_{x})<a$
for all $x\in 2^{\omega}$. Take
$K:=({2^{\omega}}\times{2^{\omega}})\setminus W.\qed$
Corollary 2.8.
Let $G\subseteq{2^{\omega}}\times{2^{\omega}}$ be a $G_{\delta}$-set such that $\mu(G|_{x})=1$ for all elements $x\in 2^{\omega}$. Then for every real number $a\in(0,1)$, there exists a compact set $K\subseteq G$ such that $\mu(K|_{x})>1-a$ for all elements $x\in 2^{\omega}$. Consequently, there exists a $\sigma$-compact set $F\subseteq G$ such that $\mu(F|_{x})=1$ for all $x\in{2^{\omega}}$.
Proof.
Let us write $G$ in the form $\bigcap_{n\in\omega}O_{n}$ where each set $O_{n}$ is open and $O_{n+1}\subseteq O_{n}$ for all natural numbers $n$. By Lemma 2.7 for each natural number $n$ there exists a compact set $K_{n}\subseteq O_{n}$ such that $\mu(K_{n}|_{x})>1-\frac{a}{2^{n+1}}$ for all $x\in 2^{\omega}$. Take $K:=\bigcap_{n\in\omega}K_{n}$. ∎
Corollary 2.9.
Let $G\subseteq{2^{\omega}}\times{2^{\omega}}$ be a $G_{\delta}$-set such that $\mu(G|_{x})=1$ for all elements $x\in 2^{\omega}$ and $\mu(\{\,x\in{2^{\omega}}:G|_{x}={2^{\omega}}\,\})=1$. Then there exists a $\sigma$-compact set $F\subseteq G$ such that $\mu(F|_{x})=1$ for all elements $x\in{2^{\omega}}$ and $\mu(\{\,x\in{2^{\omega}}:F|_{x}={2^{\omega}}\,\})=1$.
Proof.
Let $F^{\prime}\subseteq G$ be a $\sigma$-compact set from Corollary 2.8. There is an $F_{\sigma}$-set $A\subseteq\{\,x\in{2^{\omega}}:G|_{x}={2^{\omega}}\,\}$ with $\mu(A)=1$. Take $F:=F^{\prime}\cup(A\times{2^{\omega}})$. ∎
The following topological characterization of Hurewicz sets due to Just–Miller–Scheepers–Szeptycki plays a crucial role in our considerations.
Lemma 2.10 ([7, Theorem 5.7.]).
A subset $X$ of a compact space $K$ is Hurewicz if and only if for any $G_{\delta}$-set $G\subseteq K$ containing $X$, there is an $F_{\sigma}$-set set $F\subseteq K$ such that $X\subseteq F\subseteq G$.
Proof of Theorem 2.1.
Let $\mathcal{G}$ be the family of all $G_{\delta}$-sets $G$ in ${2^{\omega}}\times{2^{\omega}}$ such that $\mu(G|_{x})=1$ for all $x\in{2^{\omega}}$ and $\mu(\{\,x\in{2^{\omega}}:G|_{x}={2^{\omega}}\,\})=1$, $\mathcal{H}$ be the family of all other $G_{\delta}$-sets in ${2^{\omega}}\times{2^{\omega}}$ and let $\mathcal{F}$ be the family of all $F_{\sigma}$-sets $F$ in ${2^{\omega}}\times{2^{\omega}}$ such that $\mu(F|_{x})=1$ for all $x\in{2^{\omega}}$ and $\mu(\{\,x\in{2^{\omega}}:F|_{x}={2^{\omega}}\,\})=1$. Suppose that $\{\,(G_{\alpha},H_{\alpha}):\alpha<\mathfrak{c}\,\}$ is an enumeration of sets in $\mathcal{G}\times\mathcal{H}$ and $2^{\omega}=\{x_{\alpha}:\alpha<\mathfrak{c}\}$. By recursion on $\alpha<\mathfrak{c}$, we construct sets $F_{\alpha}\in\mathcal{F}$ and elements $u_{\alpha},w_{\alpha}\in{2^{\omega}}\times{2^{\omega}}$ with the following properties:
  1. (i)
    $\{\,u_{\beta},w_{\beta}:\beta<\alpha\,\}\cap G_{\alpha}\subseteq F_{\alpha}\subseteq G_{\alpha}$,
  2. (ii)
    $u_{\alpha},w_{\alpha}\in\bigcap\{\,F_{\beta}:\beta\leq\alpha\,\}$,
  3. (iii)
    $w_{\alpha}\notin H_{\alpha}$,
  4. (iv)
    $\pi(u_{\alpha})=x_{\alpha}$.
By Corollary 2.9 there is a set $F_{0}\in\mathcal{F}$ with $F_{0}\subseteq G_{0}$. Since $H_{0}\notin\mathcal{G}$, we have
$\mu(\{\,x\in{2^{\omega}}:F_{0}|_{x}={2^{\omega}}\,\})=1\text{\quad and \quad}\mu(\{\,x\in{2^{\omega}}:H_{0}|_{x}\neq{2^{\omega}}\,\})>0,$
or there is an element $x\in{2^{\omega}}$ such that
$\mu(F_{0}|_{x})=1\text{\quad and \quad}\mu({2^{\omega}}\setminus H_{0}|_{x})>0.$
In both cases, there is an element $x\in{2^{\omega}}$ such that
$\bigl{(}F_{0}|_{x}\bigr{)}\setminus\bigl{(}H_{0}|_{x}\bigr{)}\neq\emptyset.$
Pick elements
$u_{0}\in F_{0}\cap\bigl{(}\{x_{0}\}\times{2^{\omega}}\bigr{)}\text{\quad and\quad}w_{0}\in F_{0}\cap\bigl{(}\{x\}\times{2^{\omega}}\bigr{)}\setminus H_{0}.$
Fix an ordinal number $\alpha<\mathfrak{c}$ and assume that the objects mentioned above have been constructed for all $\beta<\alpha$. The set $\{\,u_{\beta},w_{\beta}:\beta<\alpha\,\}\cap G_{\alpha}$ has cardinality smaller than $\mathfrak{b}$, which is the critical cardinal number for the Hurewicz property, and thus this set is Hurewicz. By Lemma 2.10 and Corollary 2.9 there is a set $F_{\alpha}\in\mathcal{F}$ such that
$\{\,u_{\beta},w_{\beta}:\beta<\alpha\,\}\cap G_{\alpha}\subseteq F_{\alpha}\subseteq G_{\alpha}.$
The intersection of less than $\operatorname{cov}(\mathcal{N})$ subsets of ${2^{\omega}}$, where each subset has measure $1$, and a Borel set with positive measure, is nonempty. Since $H_{\alpha}\notin\mathcal{G}$ for all $\beta\leq\alpha$ we have
$\mu(\{\,x\in{2^{\omega}}:F_{\beta}|_{x}={2^{\omega}}\,\})=1\text{\quad and \quad}\mu(\{\,x\in{2^{\omega}}:H_{\alpha}|_{x}\neq{2^{\omega}}\,\})>0$
or there is an element $x\in{2^{\omega}}$ such that for all $\beta\leq\alpha$ we have
$\mu(F_{\beta}|_{x})=1\text{\quad and \quad}\mu({2^{\omega}}\setminus H_{\alpha}|_{x})>0.$
Since $\alpha<\operatorname{cov}(\mathcal{N})$, in both cases there is an element $x\in{2^{\omega}}$ such that
$\bigl{(}\bigcap_{\beta\leq\alpha}F_{\beta}|_{x}\bigr{)}\setminus\bigl{(}H_{\alpha}|_{x}\bigr{)}\neq\emptyset.$
Pick elements
$u_{\alpha}\in\bigcap_{\beta\leq\alpha}F_{\beta}\cap\bigl{(}\{x_{\alpha}\}\times{2^{\omega}}\bigr{)}\text{\quad and\quad}w_{\alpha}\in\Bigl{(}\bigcap_{\beta\leq\alpha}F_{\beta}\cap\bigl{(}\{x\}\times{2^{\omega}}\bigr{)}\Bigr{)}\setminus H_{\alpha}.$
This completes our recursive construction. Let $X:=\{\,u_{\alpha},w_{\alpha}:\alpha<\mathfrak{c}\,\}$.
(1) Use a characterization from Lemma 2.10. Let $G\subseteq{2^{\omega}}\times{2^{\omega}}$ be a $G_{\delta}$-set containing $X$. By (iii), none of the sets $H_{\alpha}$ contains $X$. Thus, there is an ordinal number $\alpha<\mathfrak{c}$ such that $G=G_{\alpha}$. By (i) and (ii), we have $X\subseteq F_{\alpha}\subseteq G_{\alpha}$.
(2) Let $A$ be a countable subset of ${2^{\omega}}\times{2^{\omega}}$. Then there is an ordinal number $\xi$ such that $G_{\xi}=({2^{\omega}}\times{2^{\omega}})\setminus A$. By (1) and (2), we have
$X\setminus A=X\cap G_{\xi}\subseteq F_{\xi}\subseteq G_{\xi}.$
Thus, the set $A$ is a $G_{\delta}$-set in the relative topology of $X\cup A$.
(3) Assume that $X$ contains a perfect set $P$. Then there is a countable set $A\subseteq P$ such that the set $P\setminus A$ is homeomorphic with the Baire space ${\omega^{\omega}}$. Since the set $X$ is a $\lambda^{\prime}$-set, the set $P\setminus A$ is an $F_{\sigma}$-subset of $X$. Since the Hurewicz property is preserved by $F_{\sigma}$-subsets, the set $P\setminus A$ is Hurewicz too, which is a contradiction.
(4) It follows from (iv).
(5) Apply Lemma 2.6. ∎
Corollary 2.11.
Assume that $\operatorname{cov}(\mathcal{N})=\mathfrak{b}=\mathfrak{c}$. There is a Hurewicz totally imperfect set which is not ${\mathsf{S}_{1}(\Gamma,\mathrm{O})}$.
Remark 2.12.
The proof of Theorem 2.1 actually shows that for the minimal cardinality $\kappa$ of a subset of ${2^{\omega}}$ which cannot be covered by an $F_{\sigma}$ measure zero set and any set $Y\subseteq{2^{\omega}}\times{2^{\omega}}$ with $\left|Y\right|<\kappa$, the set $Y$ is a $G_{\delta}$-subset of $Y\cup X$.
3. Menger sets
Let $\operatorname{cov}(\mathcal{M})$ be the minimal cardinality of a family of meager subsets of ${2^{\omega}}$ which covers ${2^{\omega}}$.
Theorem 3.1.
Assume that $\operatorname{cov}(\mathcal{M})=\mathfrak{c}$. Then there is a Menger totally imperfect set $X\subseteq{2^{\omega}}\times{2^{\omega}}$ with $\pi[X]={2^{\omega}}$.
Lemma 3.2.
Let $M_{0},\allowbreak M_{1},\allowbreak\dotsc$ be an increasing sequence of closed sets in ${2^{\omega}}\times{2^{\omega}}$ such that the sets $M_{n}|_{x}$ are nowhere dense in ${2^{\omega}}$ for all natural numbers $n$ and every $x\in{2^{\omega}}$. Then there is a $\sigma$-compact set $D\subseteq{2^{\omega}}\times{2^{\omega}}$ disjoint from $\bigcup_{n\in\omega}M_{n}$ such that the sets $D|_{x}$ are dense in ${2^{\omega}}$ for all elements $x\in{2^{\omega}}$.
Proof.
Let $B_{0},\allowbreak B_{1},\allowbreak\dotsc$ be an enumeration of all basic clopen sets in ${2^{\omega}}\times{2^{\omega}}$. Fix a natural number $m$. By Lemma 2.5, there is a compact set $D_{m}\subseteq B_{m}$ disjoint from the set $\bigcup_{n\in\omega}M_{n}$ such that $\pi[D_{m}]=\pi[B_{m}]$. Take $D:=\bigcup_{m\in\omega}D_{m}$. ∎
Let $\mathcal{G}$ be the family of all $G_{\delta}$-sets $G\subseteq{2^{\omega}}\times{2^{\omega}}$ such that the sets $G|_{x}$ are dense in ${2^{\omega}}$ for all elements $x\in{2^{\omega}}$.
Lemma 3.3.
Let $X\subseteq{2^{\omega}}\times{2^{\omega}}$ be a Menger set and $\mathcal{U}_{0},\allowbreak\mathcal{U}_{1},\allowbreak\dotsc$ be a sequence of open covers of $X$, where each family $\mathcal{U}_{n}$ is open in ${2^{\omega}}\times{2^{\omega}}$ and the set $\bigcap_{n\in\omega}\bigcup\mathcal{U}_{n}$ is in $\mathcal{G}$. Then there are finite sets $\mathcal{F}_{0}\subseteq\mathcal{U}_{0},\mathcal{F}_{1}\subseteq\mathcal{U}_{1},\dotsc$ such that the family $\bigcup_{n\in\omega}\mathcal{F}_{n}$ covers $X$ and the set $\bigcup_{n\in\omega}\bigcup\mathcal{F}_{n}$ is in $\mathcal{G}$.
Proof.
Let $M_{n}:={2^{\omega}}\times{2^{\omega}}\setminus\bigcup\mathcal{U}_{n}$ for all natural numbers $n$. Let $D\subseteq{2^{\omega}}$ be a $\sigma$-compact set from Lemma 3.2 applied to the sequence $M_{0},\allowbreak M_{1},\allowbreak\dotsc$. Since the set $X\cup D$ is Menger, there are finite sets $\mathcal{F}_{n}\subseteq\mathcal{U}_{n}$ for all natural numbers $n$ such that the family $\bigcup_{n\in\omega}\mathcal{F}_{n}$ covers $X\cup D$. ∎
Proof of Theorem 3.1.
Let ${\bigl{\{}\,(\mathcal{U}_{0}^{(\alpha)},\mathcal{U}_{1}^{(\alpha)},\dotsc):\alpha<\mathfrak{c}\,\bigr{\}}}$ be an enumeration of all countable sequences $\mathcal{U}_{0},\allowbreak\mathcal{U}_{1},\allowbreak\dotsc$ of open families in ${2^{\omega}}\times{2^{\omega}}$ such that the sets $\bigcap_{n\in\omega}\bigcup\mathcal{U}_{n}$ are in $\mathcal{G}$. Let ${\bigl{\{}\,(\mathcal{V}_{0}^{(\alpha)},\mathcal{V}_{1}^{(\alpha)},\dotsc):\alpha<\mathfrak{c}\,\bigr{\}}}$ be an enumeration of all other countable sequences of open families in ${2^{\omega}}\times{2^{\omega}}$, let $\{\,P_{\alpha}:\alpha<\mathfrak{c}\,\}$ be an enumeration of all perfect sets in ${2^{\omega}}\times{2^{\omega}}$ and ${2^{\omega}}=\{\,x_{\alpha}:\alpha<\mathfrak{c}\,\}$.
By recursion on $\alpha<\mathfrak{c}$, we construct finite sets $\mathcal{F}_{0}^{(\alpha)}\subseteq\mathcal{U}_{0}^{(\alpha)},\mathcal{F}_{1}^{(\alpha)}\subseteq\mathcal{U}_{1}^{(\alpha)},\dotsc$, elements $u_{\alpha},w_{\alpha}\in{2^{\omega}}\times{2^{\omega}}$ and $p_{\alpha}\in P_{\alpha}$ with the following properties:
  1. (i)
    the set $F_{\alpha}:=\bigcup_{n\in\omega}\bigcup\mathcal{F}_{n}^{(\alpha)}$ is in $\mathcal{G}$,
  2. (ii)
    $\{\,u_{\beta},w_{\beta}:\beta<\alpha\,\}\cap G_{\alpha}\subseteq F_{\alpha}$, where $G_{\alpha}:=\bigcap_{n\in\omega}\bigcup\mathcal{U}^{(\alpha)}_{n}$ ,
  3. (iii)
    $u_{\alpha},w_{\alpha}\in\bigcap_{\beta\leq\alpha}F_{\beta}$,
  4. (iv)
    $u_{\alpha},w_{\alpha}\notin\{\,p_{\beta}:\beta\leq\alpha\,\}$,
  5. (v)
    $w_{\alpha}\notin H_{\alpha}$, where $H_{\alpha}:=\bigcap_{n\in\omega}\bigcup\mathcal{V}_{n}^{(\alpha)}$,
  6. (vi)
    $\pi(w_{\alpha})=x_{\alpha}$.
Pick $p_{0}\in P_{0}$. By Lemma 3.3 there are finite sets $\mathcal{F}_{0}^{(0)}\subseteq\mathcal{U}_{0}^{(0)},\mathcal{F}_{1}^{(0)}\subseteq\mathcal{U}_{1}^{(0)},\dotsc$ such that the set $F_{0}:=\bigcup_{n\in\omega}\bigcup\mathcal{F}_{n}^{(0)}$ is in $\mathcal{G}$. The sets $F_{0}|_{x}$ are dense in ${2^{\omega}}$ for all elements $x\in{2^{\omega}}$. Since the set $H_{0}:=\bigcap_{n\in\omega}\bigcup\mathcal{V}_{n}^{(0)}$ is not in $\mathcal{G}$, there is an element $x\in{2^{\omega}}$ such that the set $H_{0}|_{x}$ is not dense in ${2^{\omega}}$, and thus
$\bigl{(}F_{0}|_{x}\bigr{)}\setminus\bigl{(}H_{0}|_{x}\bigr{)}\neq\emptyset.$
Pick elements
$u_{0}\in\bigl{(}F_{0}\cap(\{x_{0}\}\times{2^{\omega}})\bigr{)}\setminus\{p_{0}\}\text{\quad and\quad}w_{0}\in F_{0}\setminus\bigl{(}H_{0}\cup\{p_{0}\}\bigr{)}.$
Fix an ordinal number $\alpha<\mathfrak{c}$ and assume that the objects mentioned above have been constructed for all $\beta<\alpha$. Pick $p_{\alpha}\in P_{\alpha}\setminus\{\,u_{\beta},w_{\beta}:\beta<\alpha\,\}$. Since $\alpha<\operatorname{cov}(\mathcal{M})\leq\mathfrak{d}$, the set $\{\,u_{\beta},w_{\beta}:\beta<\alpha\,\}\cap G_{\alpha}$ has cardinality smaller than $\mathfrak{d}$, which is the critical cardinal number for the Menger property, and thus this set is Menger. By Lemma 3.3 there are finite sets $\mathcal{F}_{0}^{(\alpha)}\subseteq\mathcal{U}_{0}^{(\alpha)},\mathcal{F}_{1}^{(\alpha)}\subseteq\mathcal{U}_{1}^{(\alpha)},\dotsc$ such that the set $F_{\alpha}:=\bigcup_{n\in\omega}\bigcup\mathcal{F}_{n}^{(\alpha)}$ is in $\mathcal{G}$ and
$\{\,u_{\beta},w_{\beta}:\beta<\alpha\,\}\cap G_{\alpha}\subseteq F_{\alpha}.$
The intersection of less than $\operatorname{cov}(\mathcal{M})$ dense $G_{\delta}$-sets in ${2^{\omega}}$ and a subset of ${2^{\omega}}$ with nonempty interior is nonempty. For each $\beta\leq\alpha$ the sets $F_{\beta}|_{x}$ are dense $G_{\delta}$-sets in ${2^{\omega}}$ for all elements $x\in{2^{\omega}}$. Since the set $H_{\alpha}:=\bigcap_{n\in\omega}\bigcup\mathcal{V}^{(\alpha)}_{n}$ is not in $\mathcal{G}$, there is $x\in{2^{\omega}}$ such that the set $H_{\alpha}|_{x}$ is not dense in ${2^{\omega}}$, i.e., the set ${2^{\omega}}\setminus H_{\alpha}|_{x}$ has nonempty interior. Since $\alpha<\operatorname{cov}(\mathcal{M})$, we can pick elements
$u_{\alpha}\in\bigl{(}\bigcap_{\beta\leq\alpha}F_{\beta}\cap(\{x_{\alpha}\}\times{2^{\omega}})\bigr{)}\setminus\{\,p_{\beta}:\beta\leq\alpha\,\}$
and
$w_{\alpha}\in\bigcap_{\beta\leq\alpha}F_{\beta}\setminus\bigl{(}H_{0}\cup\{\,p_{\beta}:\beta\leq\alpha\,\}\bigr{)}$
This completes our recursive construction. Let $X:=\{\,u_{\alpha},w_{\alpha}:\alpha<\mathfrak{c}\,\}$.
Let $\mathcal{U}_{0},\allowbreak\mathcal{U}_{1},\allowbreak\dotsc$ be a sequence of open covers of $X$. For each ordinal number $\alpha<\mathfrak{c}$ we have $w_{\alpha}\notin H_{\alpha}=\bigcap_{n\in\omega}\bigcup\mathcal{V}_{n}^{(\alpha)}$, and thus none of the sequences $\mathcal{V}_{0}^{(\alpha)},\mathcal{V}_{1}^{(\alpha)},\dotsc$ is a sequence of covers of $X$. Consequently, there is an ordinal number $\alpha<\mathfrak{c}$ such that the sequence $\mathcal{U}_{0},\allowbreak\mathcal{U}_{1},\allowbreak\dotsc$ is equal to $\mathcal{U}_{0}^{(\alpha)},\mathcal{U}_{1}^{(\alpha)},\dotsc$ By (ii) we have
$\{\,u_{\beta},w_{\beta}:\beta<\alpha\,\}=\{\,u_{\beta},w_{\beta}:\beta<\alpha\,\}\cap G_{\alpha}\subseteq F_{\alpha}=\bigcup_{n\in\omega}\bigcup\mathcal{F}_{n}^{(\alpha)}$
and by (iii) we have
$\{\,u_{\beta},w_{\beta}:\alpha\leq\beta<\mathfrak{c}\,\}\subseteq F_{\alpha}=\bigcup_{n\in\omega}\bigcup\mathcal{F}_{n}^{(\alpha)}.$
Then the family $\bigcup_{n\in\omega}\mathcal{F}_{n}^{(\alpha)}$ covers $X$, and thus the set $X$ is Menger.
By (iv) the set $X$ is totally imperfect and by (v) we have $\pi[X]={2^{\omega}}$. ∎
Corollary 3.4.
Assume that $\operatorname{cov}(\mathcal{M})=\mathfrak{c}$. There is a Menger totally imperfect set which is not ${\mathsf{S}_{1}(\Gamma,\mathrm{O})}$.
4. More consequences
4.1. On the problem of Nowik and Tsaban
The set from Theorem 2.1 provides a negative answer to the question from Problem 1.2 under $\operatorname{cov}(\mathcal{N})=\mathfrak{b}=\mathfrak{c}$. On the other hand it is consistent with ZFC that all Hurewicz totally imperfect sets have cardinalities strictly smaller than $\mathfrak{c}$. By the results of Miller [10] and Zakrzewski [27], this holds in the Cohen model, and by the result of Miller [10] also in the Sacks model. In these cases an affirmative answer to Problem 1.2 follows from Theorem 1.1.
4.2. Everywhere meager and everywhere null
A subset $A$ of a Polish space $X$ is universally meager [27, 28, 3, 4] if for any Borel isomorphism between $X$ and a Polish space $Y$, the image of $A$ under the map is meager. A set $X\subseteq{2^{\omega}}$ is everywhere meager [8] if for each set $a\in{[\omega]^{\omega}}$ the set $X_{a}:=\{\,x\upharpoonright a:x\in X\,\}$ is meager in $2^{a}$. Kraszewski proved that there is a universally meager set in ${2^{\omega}}$ which is not everywhere meager [8, Theorem 4.1]. We give a consistently negative answer to a weaker version of the following question.
Problem 4.1 (Kraszewski [8, Problem 1]).
Is every PMT subset of ${2^{\omega}}$ everywhere meager?
A set $X\subseteq{2^{\omega}}$ is countably perfectly meager if for any sequence $P_{0},\allowbreak P_{1},\allowbreak\dotsc$ of perfect subsets of ${2^{\omega}}$, there is an $F_{\sigma}$-set $F\subseteq{2^{\omega}}$ containing $X$ and the sets $P_{n}\cap F$ are meager in $P_{n}$ for all natural numbers $n$. Equivalently, the set $X$ is countably perfectly meager if for each perfect set $P$ and a countable set $Q$ there is an $F_{\sigma}$-set $F$ containing $X$ such that for every $q\in Q$, the set $F\cap(q+P)$ is meager in $q+P$. Every Hurewicz totally imperfect subset of ${2^{\omega}}$ is countably perfectly meager [15]. By the results of Pol–Zakrzewski [15] and Weiss–Zakrzewski [26] there is consistently a universally meager set which is not countably perfectly meager.
Lemma 4.2.
If $X\subseteq{2^{\omega}}$ is a set whose continuous image into ${2^{\omega}}$ is not meager, then a homeomorphic copy of $X$ in ${2^{\omega}}$ is not everywhere meager.
Proof.
Let $f\colon X\to{2^{\omega}}$ be a continuous map, whose image $f[X]$ is not meager in ${2^{\omega}}$. Then the set
$Y:=\{\,(x,f(x)):x\in X\,\},$
the graph of the function $f$, is a homeomorphic copy of $X$ in ${2^{\omega}}\times{2^{\omega}}$. Fix disjoint sets $a,b\in{[\omega]^{\omega}}$ with $a\cup b=\omega$. Let $\psi_{a}\colon{2^{\omega}}\to 2^{a}$ and $\psi_{b}\colon{2^{\omega}}\to 2^{b}$ be homeomorphisms. Then the map $\psi:=\psi_{a}\times\psi_{b}$ is a homeomorphism from ${2^{\omega}}\times{2^{\omega}}$ to $2^{a}\times 2^{b}$. Let $\phi\colon 2^{a}\times 2^{b}\to{2^{\omega}}$ be a homeomorphism such that $\phi(x,y):=x\cup y$ for all $(x,y)\in 2^{a}\times 2^{b}$. Then the set
$Z:=(\phi\circ\psi)[Y]$
is a homeomorphic copy of $X$ such that the set $Z_{b}=\psi_{b}[f[X]]$ is not meager in $2^{b}$. ∎
The following result answers Problem 4.1 in which PMT sets are replaced by countably perfectly meager sets in the negative.
Corollary 4.3.
Assume that $\operatorname{cof}(\mathcal{N})=\mathfrak{b}=\mathfrak{c}$. Then there is a totally imperfect Hurewicz (and thus countably perfectly meager) subset of ${2^{\omega}}$ which is not everywhere meager.
Proof.
Let $X$ be a set from Theorem 2.1. By Lemma 4.2, there is a homeomorphic copy $Y$ of $X$ in ${2^{\omega}}$ which is not everywhere meager. Since $Y$ is totally imperfect and Hurewicz, by the result of Pol–Zakrzewski [15], it is countably perfectly meager. ∎
According to a characterization of universally meager sets due to Zakrzewski [27, Theorem 2.1], a subset of a Polish space is universally meager if and only if every Borel isomorphic image of this set into a Polish space is meager. It follows that universally meager sets are preserved by homeomorphisms. By the result of Pol and Zakrzewski [15, Theorem 3.1.3] consistently countably perfectly meager sets are not preserved by homeomorphisms. We do not know whether of everywhere meager sets are preserved by homeomorphisms.
4.3. The property $\mathsf{U}_{n}(\Gamma,\Gamma)$
A set $X$ satisfies $\mathsf{U}_{n}(\Gamma,\Gamma)$ if for any sequence $\mathcal{U}_{0},\allowbreak\mathcal{U}_{1},\allowbreak\dotsc$ of open $\gamma$-covers of $X$, there are sets $\mathcal{F}_{0}\subseteq\mathcal{U}_{0},\mathcal{F}_{1}\subseteq\mathcal{U}_{1},\dotsc$ such that $\left|\mathcal{F}_{n}\right|=n$ for all $n$’s and the family $\{\,\bigcup\mathcal{F}_{n}:n\in\omega\,\}$ is a $\gamma$-cover of $X$. This property was investigated by Tsaban [23] in the context of sets with the Hurewicz property and the property ${\mathsf{S}_{1}(\Gamma,\Gamma)}$. A set $X$ satisfies ${\mathsf{S}_{1}(\Gamma,\Gamma)}$ if for any sequence $\mathcal{U}_{0},\allowbreak\mathcal{U}_{1},\allowbreak\dotsc$ of open $\gamma$-covers of $X$, there are sets $U_{0}\in\mathcal{U}_{0},U_{1}\in\mathcal{U}_{1},\dotsc$ such that the family $\{\,U_{n}:n\in\mathbb{N}\,\}$ is a $\gamma$-cover of $X$. Each set satisfying ${\mathsf{S}_{1}(\Gamma,\mathrm{O})}$ is totally imperfect [23, Lemma 1.16]. Using a similar argument, we can show that any set satisfying $\mathsf{U}_{n}(\Gamma,\Gamma)$ is totally imperfect, too. A $\mathfrak{b}$-scale set is a subset of the Cantor space ${\operatorname{P}(\omega)}$ of the form $\{\,x_{\alpha}:\alpha<\mathfrak{b}\,\}\cup\mathrm{Fin}$, where $\{\,x_{\alpha}:\alpha<\mathfrak{b}\,\}$ is an unbounded subset of ${[\omega]^{\omega}}$ such that for every $\alpha<\beta<\mathfrak{b}$, we have $x_{\alpha}\leq^{*}x_{\beta}$. Such a set exists in ZFC and is an example of a Hurewicz totally imperfect set given by Bartoszyński–Shelah [5]. The only known example which distinguishes the properties ${\mathsf{S}_{1}(\Gamma,\Gamma)}$ and Hurewicz in the class of totally imperfect sets is a $\mathfrak{b}$-scale set [16] under $\mathfrak{b}=\mathfrak{c}$. By the results of Tsaban [23, Theorems 3.3 and 3.8] any $\mathfrak{b}$-scale set satisfies $\mathsf{U}_{n}(\Gamma,\Gamma)$, and thus any continuous image of a $\mathfrak{b}$-scale set into ${2^{\omega}}$ is totally imperfect. Tsaban asked the following question.
Problem 4.4 ([23, Problem 3.9]).
Is there a totally imperfect Hurewicz set which is not $\mathsf{U}_{n}(\Gamma,\Gamma)$?
Our example constructed in Theorem 2.1 provides a negative answer to the above question under $\operatorname{cov}(\mathcal{N})=\mathfrak{b}=\mathfrak{c}$.
5. Comments and open problems
Problem 5.1.
Does the existence of a totally imperfect Hurewicz set of size $\mathfrak{c}$ imply that there is a totally imperfect Hurewicz set which is not PMT?
Problem 5.2.
Let $X\subseteq{2^{\omega}}$. Are the following statements equivalent?
  1. (1)
    The set $X$ is Hurewicz and totally imperfect which can be mapped continuously onto ${2^{\omega}}$.
  2. (2)
    The set $X$ is not PMT?
Problem 5.3.
Is there a Menger set whose all continuous images into ${2^{\omega}}$ are totally imperfect and which does not satisfy ${\mathsf{S}_{1}(\Gamma,\mathrm{O})}$?